Part A. For the given system, determine the best method (Graphing, Substitution, Addition) for solving.
3x - 2y = 8
x = -2y
Part B. Explain why you chose that method.
Part C. Solve the system. Give the final solution expressed as (x,y) point. (Showing your steps

Answer by stanbon(75887)

For the given system, determine the best method (Graphing, Substitution, Addition) for solving.
3x - 2y = 8
x = -2y
Part B. Explain why you chose that method.
Substitution because the 2nd equation is all set up for substitution.
------
Part C. Solve the system. Give the final solution expressed as (x,y) point. (Showing your steps
---
Substitute for "x" and solve for "y":
3(-2y) - (2y) = 8
----
-6y-2y = 8
-8y = 8
y = -1
---
Solve for "x":
x = -2y
x = -2*-1 = 2
=====================
solution: (2,-1)
=======================
cheers,
Stan H.
===============